# onlineReferences

### v5

@@ 2.5   Comparing DID-IV to Fuzzy DID
The main difference between dCDH and this article lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.121212de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. Therefore, we rewrite their target parameter by the notation used in this article.
@@
$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
where the switcher (SW) is the units who become treated in period $1$.
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). Because the target parameter is different from each other, the identifying assumptions in Fuzzy DID designs are also different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In online Appendix Section E, we first show that the identifying assumptions in Fuzzy DID designs ex ante impose strong restrictions on treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups. Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 3-4, which are in line with Tables 1-2.131313Because dCDH consider repeated cross section settings, we can observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.
Comparing Tables 1-2 with Tables 3-4, one can figure out the difference between dCDH’s restrictions and ours. In our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption. By contrast, dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
@@
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by dCDH’s identifying assumptions.
Next, under Tables 3-4, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers. The other causal parameter measures the treatment effects among the time compliers $CM^{T}$. [^14] Intuitively, we cannot identify this parameter in DID-IV settings because the untreated potential outcomes in period $t=1$ among the time compliers are not observed in both groups (see Tables 3-4). This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher, as this parameter is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by the instrument.
Finally, we revisit dCDH’s argument regarding the use of the Wald-DID estimand, and show that this problem is less serious than one might expect, and does not arise once a researcher imposes Assumption 6 instead of the parallel trends assumption in the untreated outcome.
@@
First, we show that their argument is based on the parallel trends assumption in DID designs. Recall that in our DID-IV design, we do not make the parallel trends assumption in the untreated outcome, but make the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
@@
Next, we argue that the different type of the parallel trends assumption, Assumption 6, is suitable for DID-IV settings, and the parallel trends assumption in the untreated outcome would be difficult to justify in practice for three reasons. First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Duflo (2001), this assumption requires the data to contain units with zero education attainment, which is unrealistic in practice. Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.
@@ 3.1   Set up
For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
Assumption 8 requires that once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 8, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$.
@@
$D_{i,t}=D_{i,t}^{\infty}+\sum_{1\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
@@ 3.2   Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$. [^16] Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher. For example, in a job training program setting considered in Subsection 2.5, each CLATT would be of particular interest if the instrument $Z_{i,t}$ represents an important policy lever for the researchers (e.g., nudge); the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
@@ 3.3   Identification assumptions in staggered DID-IV designs
In this subsection, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
@@
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d)\hskip 8.53581pta.s.$
@@
$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{S}(E_{i})\hskip 5.69054pt\text{and for all}\hskip 5.69054ptl\geq 0.$
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 11 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
@@
[…]
@@
Assumption 13 (Parallel trends assumption in the treatment in multiple time periods).
[…]
@@
Assumption 13 is a parallel trends assumption in the treatment in multiple periods and cohorts. It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument. Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
Assumption 14 (Parallel trends assumption in the outcome in multiple time periods).
[…]
Assumption 14 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts. It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned. From the discussion in Subsection 2.4, this assumption can be interpreted as requiring the same expected time gain across cohorts and over time; that is, the effects of time on the outcome through treatment are the same on average across cohorts and over time.
Assumption 15 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]>0.$
Assumption 15 is a relevance condition in multiple period settings. With a binary treatment, Assumption 11 and 15 implies that the fraction of the compliers should be non-negative for all $e$ and $l\geq 0$:
$E[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]=Pr(CM_{e,l}|E_{i}=e)>0.$
This condition ensures that each $CLATT_{e,l}$ is well defined.

### v6

@@ 2.4  Comparing DID-IV with Fuzzy DID
The main difference between dCDH and this paper lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
@@
$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers, SW).
Because the target parameters differ, the identifying assumptions underlying Fuzzy DID designs also differ from those in DID-IV designs, though there are similarities as well. [^13] Specifically, dCDH also impose Assumptions 1-4 and 7 in this paper. The main differences are as follows. First, dCDH assume the stable treatment rate assumption in an unexposed group (see Assumption 22 in dCDH), whereas we assume the parallel trends assumption in the treatment (Assumption 5). Second, dCDH assume the treatment participation equation (see Assumption 33 in dCDH), which imposes the monotonicity with respect to time $T$ in addition to the monotonicity with respect to instrument $Z$ (which corresponds to Assumption 3 in this paper). Finally, while we impose a parallel trends (PT) assumption on unexposed outcomes, dCDH impose a PT assumption on untreated outcomes (see Assumption $4$ in dCDH).
In Online Appendix D, we first examine how these differences imply different restrictions on treatment adoption behavior across the two designs. We begin by describing the heterogeneity in treatment adoption behavior under the setting considered here. Specifically, we introduce the group variable $G^{T}=(D_{0}(0),D_{1}(0))$, which characterizes treatment adoption over time when the instrument is $Z=0$ in the second period. We define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers. Using the group variables $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case) and $G^{T}$, we then partition units into eight types within each group, as summarized in Tables 1-2.
Next, we investigate which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively. Specifically, Tables 12 and Tables 34 summarize all latent treatment adoption types under DID-IV and Fuzzy DID respectively, where the types painted in gray color are excluded by the identifying assumptions of each design. Here, in Table 3, we exclude the $DF^{Z}$ and the $DF^{T}$ by the monotonicity assumptions with respect to instrument $Z$ and time $T$, which are implied by the treatment participation equation (see Lemma 1 in Online Appendix D.2.3).
@@
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by the monotonicity assumption with respect to instrument $Z$.
Table 3: Exposed group ($z=1$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 4: Unexposed group ($z=0$)
observed counterfactual
[…]
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by the identifying assumptions in dCDH.
By comparing Tables 12 with Tables 34, we obtain two implications. First, the restrictions imposed by Fuzzy DID designs are stronger than those imposed by DID-IV designs. Second, while the restrictions under DID-IV designs are symmetric across exposed and unexposed groups, those under Fuzzy DID designs are asymmetric.
Building on Tables 34, we next clarify the difference in target parameter between the two designs. Specifically, we show that dCDH’s target parameter, the SLATET, can be expressed as a weighted average of two distinct causal parameters. One parameter captures the treatment effects for units of type $CM^{Z}\land NT^{T}$, a subpopulation of the compliers $CM^{Z}$, while the other captures the treatment effects among the type $AT^{Z}\land CM^{T}$ (case (i)).
Importantly, this decomposition depends on the direction of the two monotonicity assumptions. If $DF^{Z}$ and $CM^{T}$ are excluded by the two monotonicity conditions in dCDH, the SLATET identifies the treatment effects for units of type $CM^{Z}\land NT^{T}$ (case (ii)). If $CM^{Z}$ and $DF^{T}$ are excluded, the SLATET identifies the treatment effects for units of type $AT^{Z}\land CM^{T}$ (case (iii)), as formally established in Theorem 6 in Appendix D.3.
This decomposition result has several important implications. First, empirical researchers should specify the direction of the two monotonicity conditions ex ante if they wish to understand which latent type’s causal effect is identified by the SLATET. Second, the SLATET may fail to be the policy-relevant parameter (Heckman and Vytlacil (2001)), as it is generally contaminated by treatment effect for units whose treatment status is affected by time but not affected by the instrument (see cases (i) and (iii)). Finally, even when the two monotonicity conditions ensure that the SLATET is policy relevant, it identifies the treatment effects for a narrower population than the LATET (see case (ii)).
Finally, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs. In DID-IV, we view the Wald–DID estimand as a natural estimand for identifying the LATET. By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption to identify the SLATET.141414dCDH also show that the Wald–DID estimand additionally requires homogeneity of the SLATET between the exposed and unexposed groups in order to identify the SLATET in the exposed group when the stable treatment rate condition in the unexposed group is not satisfied. In earlier work, Blundell and Dias (2009) make a similar argument in Subsection F.2 of Section III. In Remark 4 of Section D.3 in Online Appendix D, however, we show that this homogeneity condition cannot be straightforwardly interpreted as requiring homogeneous treatment effects between the two groups. We demonstrate that this difference arises because Fuzzy DID and DID-IV rely on different types of the PT assumption. Given this, we also discuss which PT assumption is more suitable for DID-IV settings.
@@
Specifically, we first show that the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH impose the PT assumption on untreated outcomes. Recall that in DID-IV settings, units are allowed to adopt the treatment without instrument during the two periods. As a result, the PT assumption on untreated outcomes is insufficient to capture the average time trends of the outcome even in the unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. By contrast, this issue does not arise in DID-IV designs because we impose the PT assumption on unexposed outcomes, which is sufficient to capture the average time trends of the outcome for the unexposed group.
@@
Next, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings for three reasons. First, the PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes. This is because the latter relies on variation in treatment, while the former exploits variation in instrument. Second, in most applications of DID-IV methods, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Duflo (2001), the PT assumption on untreated outcomes would require the data to include units with zero educational attainment, which is unrealistic in practice. Finally, in DID-IV settings, while the PT assumption on unexposed outcomes is indirectly testable (see Section 5.3 in this paper), the PT assumption on untreated outcomes is difficult to assess using pre-exposed period data, as some units may already adopt the treatment before period 0.
We conclude this section by providing guidance for empirical researchers choosing between DID-IV and Fuzzy DID designs in practice. First, when the goal is to identify a policy-relevant parameter, DID-IV designs are more attractive. Under Fuzzy DID designs, the target parameter may not be policy relevant in general. Second, when researchers are interested in identifying the SLATET, they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed by Fuzzy DID designs. If these restrictions appear questionable, researchers may instead target the LATET and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument. Finally, if researchers wish to assess the plausibility of the PT assumption, DID-IV designs are more appealing because the PT assumption imposed under Fuzzy DID is generally difficult to test using the pre-exposed data.
@@ 3.1  Set up
$Z_{i,1}=0$ for all $i$. For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
Assumption 8 requires that no one is exposed to the instrument in time $t=1$ and once units start getting exposed to the instrument, units remain exposed to that instrument. In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g. Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
Given Assumption 8, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{2,3,\dots,T,\infty\}$. Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$. Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset. Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
@@
$D_{i,t}=D_{i,t}^{\infty}+\sum_{2\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
@@ 3.2  Target parameter in staggered DID-IV designs
Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$.
@@ 3.3  Identification assumptions in staggered DID-IV designs
In this subsection, we establish the identification assumptions in staggered DID-IV designs.
In staggered DID-IV designs, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
$w^{DID}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=\infty])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])},$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. Note that this estimand is the Wald-DID estimand, where the pre-exposed period is e1e-1 and the control group is the never exposed cohort. Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
We consider the following identification assumptions for each $w^{DID}_{e,l}$ to capture the $CLATT_{e,e+l}$.
@@
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d).$
@@
$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl\geq 0.$
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 11 requires that the individual exposed effect in the first stage should be non-negative or non-positive during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e\in\mathcal{E}$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e\in\mathcal{E}$.
@@
$D_{i,e+l}^{e}=D_{i,e+l}^{\infty}\hskip 8.53581pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl<0.$
@@
Assumption 13 (Parallel trends assumption in the treatment based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=\infty].$
@@
Assumption 13 is a parallel trends assumption in the treatment based on the never exposed cohort. It requires that, in the absence of exposure to the instrument, the average evolution of the treatment would have followed the same path between cohort $e$ and the never exposed cohort $\infty$. Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
Assumption 14 (Parallel trends assumption in the outcome based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=\infty].$
Assumption 14 is a parallel trends assumption in the outcome based on the never exposed cohort. It requires that, in the absence of exposure to the instrument, the expected evolution of the outcome under no exposure to the instrument would have been the same, on average, between cohort $e$ and the never exposed cohort $\infty$.
Assumption 15 (Relevance condition based on a never exposed cohort).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])>0.$
Assumption 15 is a relevance condition in multiple period settings, ensuring that the Wald-DID estimand $w^{DID}_{e,l}$ is well defined.
The following theorem shows that under Assumptions 815, each Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
Theorem 2.
If Assumptions 815 hold, the Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
$w^{DID}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$.
Proof.
See Appendix. ∎
In some applications, however, there may be no never exposed cohort, or its sample size may be too small to serve as a reliable control group. Moreover, the parallel trends assumptions based on the never exposed cohort may be viewed as less credible in certain empirical settings. In such cases, it is natural to instead use the not-yet-exposed cohorts as the control group. Accordingly, we consider the following alternative Wald-DID estimand:
$w^{DID,ny}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])},$
for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. In this Wald-DID estimand, the control group is the not-yet-exposed cohorts, that is, the cohorts not exposed to the instrument by time $t=e+l$.
If we adopt the above estimand $w^{DID,ny}_{e,l}$, we can replace Assumptions 1315 with Assumptions 1618 below.
Assumption 16 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|Z_{i,t}=0,E_{i,e}=0].$
Assumption 17 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|Z_{i,t}=0,E_{i,e}=0].$
Assumption 18 (Relevance condition based on not-yet-exposed cohorts).
$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])>0.$
The following theorem shows that under Assumptions 812 and 1618, each Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
Theorem 3.
If Assumptions 812 and 1618 hold, the Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
$w^{DID,ny}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
Proof.
See Appendix. ∎